# Fine-tuning can Help Detect Pretraining Data from Large Language Models

### v1

@@ 6 RELATED WORK
7 CONCLUSION
Limitations
Our method requires to collect a few examples that belong to the same domain but are not involved in the training. Generally, we can utilize the data content published after the release of the LLM. Therefore, our method is applicable for detecting benchmarks or copyrighted resources in a specific domain (e.g., math tests, magazines). The diversity of the test set may make it challenging to construct an effective auxiliary dataset of unseen data. In addition, the effectiveness of our method may reflect the distribution difference between members and non-members in existing popular benchmarks, which is inconsistent with the assumption of MIAs. It might be necessary to inspect the dataset splitting in the common benchmarks of detecting pretraining data.
Appendix A Details of Datasets
Previous works construct benchmark datasets to evaluate the performance of detection methods for pretraining data detection. Following the prior literature, we conduct experiments on 4 benchmark datasets: WikiMIA (Shi et al., 2024) selects old Wikipedia event data as member data by leveraging the Wikipedia data timestamp and the model release date, since Wikipedia is a commonly pretraining data source. BookMIA (Shi et al., 2024), which contains excerpts from copyrighted books in the Books3 subset of the Pile dataset (Gao et al., 2020), can be used for detecting potential copyright infringement in training data. ArXivTection (Duarte et al., 2024) is a curated collection of research articles sourced from arXiv. BookTection (Duarte et al., 2024), which comprises passages from 165 books, is constructed based on BookMIA.
Appendix B Experimental Detail
B.1 Dataset split
We report the performance of detection methods on WikiMIA and ArXivTection datasets across various large language models. To construct the fine-tuning dataset, we randomly split 30% of the dataset using a seed of 42 and select all non-members from this subset. The remaining 70% of the dataset is used for testing. The detailed information of the constructed dataset is shown in Table 7.
@@ B.1 Dataset split
B.2 Copyrighted book detection
To conduct experiments of copyrighted book detection on BookMIA and BookTection, we randomly split 30% of the dataset using a seed of 42 and select all non-members from this subset as the fine-tuning dataset. Subsequently, we randomly sample 500 members and non-members from the remaining 70% of the datasets, constructing a balanced validation set of 1,000 examples. The remaining data is used for testing. The detailed information dataset split is shown in Table 8.
@@ B.2 Copyrighted book detection
B.3 Fine-tuning with members
To investigate the impact of model fine-tuning with different fine-tuning datasets on pretraining data detection, we construct three kinds of fine-tuning datasets. In this ablation, we sample members (Mem) and non-members (Non) from WikiMIA as fine-tuning datasets, respectively. In addition, we randomly sample data from WikiMIA to construct a fine-tuning dataset (All). The details of fine-tuning datasets are shown in Table 9
@@ B.3 Fine-tuning with members
      Test Set       536       549       1,085
B.4 Temporal shift
We show the temporal shift between members and non-members in the WikiMIA dataset, the illustration is presented in Table 10.
Table 10: An illustration of temporal shift between the member and non-member distributions.
@@ B.4 Temporal shift
Appendix C Detailed Experimental Results
C.1 Fine-tuned Score Deviation
We show in Figure 5 the score distribution between members and non-members using various scoring functions on ArXivTection. The results also demonstrate that our FSD score significantly increases the gap between non-members and members compared to the perplexity and Min-kscores, thus enabling more effective pretraining data detection.
(a) Perplexity
(b) Min-k%
(c) FSD with Perplexity
(d) FSD with Min-k%
C.2 Detailed results of experiment
Is FSD effective with different-sized models?
Is our method affected by distribution difference?

### v2

@@ Appendix A RELATED WORK
Appendix B Details of Datasets
Previous works construct benchmark datasets to evaluate the performance of detection methods for pretraining data detection. Following the prior literature, we conduct experiments on 5 benchmark datasets: WikiMIA (Shi et al., 2024) selects old Wikipedia event data as member data by leveraging the Wikipedia data timestamp and the model release date, since Wikipedia is a commonly pretraining data source. BookMIA (Shi et al., 2024), which contains excerpts from copyrighted books in the Books3 subset of the Pile dataset (Gao et al., 2020), can be used for detecting potential copyright infringement in training data. ArXivTection (Duarte et al., 2024) is a curated collection of research articles sourced from ArXiv. BookTection (Duarte et al., 2024), which comprises passages from 165 books, is constructed based on BookMIA. We also conducted experiments on the Pile dataset (Maini et al., 2024), which is large-scale text dataset for training language models, including text data from various sources such as books, GitHub, and website content.
Appendix C Experimental Detail
C.1 Dataset split
We report the performance of detection methods on WikiMIA and ArXivTection datasets across various large language models. In our experiments, we first randomly sample 30% of the dataset, and then select all non-members from this subset to construct the fine-tuning dataset. The remaining 70% of the dataset is used for testing. The detailed information of the constructed dataset is shown in Table 8.
@@ C.1 Dataset split
C.2 Copyrighted book detection
To conduct experiments of copyrighted book detection on BookMIA and BookTection, we randomly sample 30% of the dataset and select all non-members from this subset as the fine-tuning dataset. Subsequently, we randomly sample 500 members and non-members from the remaining 70% of the datasets, constructing a balanced validation set of 1,000 examples for evaluation. The detailed information dataset split is shown in Table 9.
@@ C.2 Copyrighted book detection
C.3 Fine-tuning with members
To investigate the impact of model fine-tuning with different fine-tuning datasets on pretraining data detection, we construct three types of fine-tuning datasets. In this ablation, we sample members (Mem) and non-members (Non) from WikiMIA as fine-tuning datasets, respectively. In addition, we randomly sample data from WikiMIA to construct a fine-tuning dataset (All). The details of fine-tuning datasets are shown in Table 10
@@ C.3 Fine-tuning with members
      Test Set       599       558       1,157
C.4 Temporal shift
In Table 11, we provide some illustrations of the temporal shift between members and non-members in the WikiMIA dataset.
Table 11: Illustrations of temporal shift between the member and non-member distributions.
@@ C.4 Temporal shift
Appendix D Detailed Experimental Results
@@
D.1 Fine-tuned Score Deviation
We show in Figure 5 the score distribution between members and non-members using two scoring functions on ArXivTection. The results also demonstrate that our FSD score significantly increases the gap between non-members and members compared to the perplexity and Min-kscores, thus enabling more effective pretraining data detection.
(a) Perplexity
(b) Min-k%
(c) FSD with Perplexity
(d) FSD with Min-k%
D.2 Detailed results of experiment
Is FSD effective with different-sized models?
Is our method affected by distribution difference?
D.3 Additional results
The performance of our method on the Pile dataset
We also conduct experiments on the Pile dataset. Concretely, following prior work (Maini et al., 2024), we evaluate our methods on the twenty subsets of the Pile dataset. Here, the validation set of the pile dataset was not trained on the Pythia models (Biderman et al., 2023). Thus, we perform experiments on the Pythia-6.9B model, utilizing the training and validation sets as members and non-members, respectively. For each dataset, we randomly sample a few non-members with a sample ratio of 0.3 from the validation set for fine-tuning. Then, we evaluate our method on a balanced dataset composed of members and non-members. Notably, in our experiments, there is no overlap between the fine-tuning dataset and the evaluation data.
@@ D.3 Additional results
In Table 15, the results show that our method improves the performance of baselines on most subsets of the Pile dataset under the Pythia-6.9B model. For example, our FSD improves the AUC score of the perplexity-based method from 0.528 to 0.885 on BookC2, a significant direct improvement of 67%. At the same time, our FSD improves the average AUC score of the perplexity-based method from 0.503 to 0.625 on the pile dataset, a notable direct improvement of 24.3%. This demonstrates the effectiveness of our method in the Pile dataset.
Table 15: AUC score for pretraining data detection with baselines and our method from the Pythia-6.9B on the Pile dataset. Base and +Ours respectively denote the baseline methods and our method. Bold shows the superior result.
Method Wiki BookC2 Gutenberg HackerNews Enron
Base +Ours Base +Ours Base +Ours Base +Ours Base +Ours
Perplexity 0.471 0.614 0.528 0.885 0.528 0.661 0.471 0.565 0.510 0.678
Lowercase 0.466 0.626 0.518 0.725 0.546 0.551 0.450 0.512 0.484 0.659
Zlib 0.496 0.619 0.477 0.907 0.496 0.686 0.474 0.550 0.560 0.667
Min-k0.512 0.611 0.510 0.841 0.536 0.612 0.498 0.535 0.570 0.646
Method CC arXiv Europarl FreeLaw GitHub
Base +Ours Base +Ours Base +Ours Base +Ours Base +Ours
Perplexity 0.541 0.546 0.514 0.505 0.514 0.601 0.478 0.515 0.509 0.548
Lowercase 0.502 0.547 0.523 0.530 0.521 0.556 0.476 0.507 0.491 0.513
Zlib 0.529 0.576 0.540 0.505 0.462 0.609 0.492 0.503 0.491 0.562
Min-k0.557 0.542 0.515 0.502 0.512 0.583 0.492 0.500 0.513 0.551
Method Books3 Nih OpenWebtext2 PhilPapers OpenSubtitles
Base +Ours Base +Ours Base +Ours Base +Ours Base +Ours
Perplexity 0.560 0.509 0.463 0.599 0.490 0.580 0.571 0.869 0.525 0.521
Lowercase 0.550 0.524 0.608 0.512 0.486 0.547 0.633 0.718 0.538 0.528
Zlib 0.550 0.581 0.416 0.599 0.475 0.586 0.678 0.871 0.550 0.530
Min-k0.552 0.554 0.463 0.560 0.510 0.567 0.606 0.826 0.525 0.535
Method StackExchange Math YoutubeSubtitles USPTO Ubuntu
Base +Ours Base +Ours Base +Ours Base +Ours Base +Ours
Perplexity 0.640 0.678 0.530 0.504 0.392 0.756 0.537 0.606 0.282 0.767
Lowercase 0.579 0.641 0.508 0.513 0.495 0.546 0.510 0.582 0.496 0.476
Zlib 0.595 0.686 0.513 0.502 0.445 0.736 0.484 0.604 0.423 0.592
Min-k0.637 0.670 0.524 0.510 0.380 0.692 0.549 0.596 0.329 0.561
Fine-tuning on BookC2 with varying data size
We also conduct experiments on the BookC2 subset of the Pile dataset under the Pythia-6.9B model to investigate the effect of varying the fine-tuning data size on the pretraining data detection. Specifically, we randomly sample varying amounts of non-members (0, 50, 100, 150, 200, 250, 300, 350, 400, 450, 500) from the validation set of the BookC2 as fine-tuning datasets. In addition, we sample 1400 members and non-members from the train and validation sets of the BookC2 to construct a balanced test set of 2800 examples.
Figure 6(a) shows that our method achieves better performance as the size of the fine-tuning dataset increases. Notably, our method is highly data-efficient, achieving significant improvements with only a few non-members for fine-tuning. For instance, our method improves the AUC score of the Zlib method from 0.48 to 0.78, by leveraging only 100 non-member data for fine-tuning. In addition, the results of the TPR@5%FPR score show a similar trend, which can be found in Figure 6(b).
Fine-tuning using non-members from different domains
Our method requires a few non-member data from a specific domain for fine-tuning. This raises a question: how does our method perform when fine-tuned on non-member data from a different domain? To investigate the performance of our method when fine-tuning using data from an unrelated domain. Firstly, we randomly sample 231 and 238 non-members from the WikiMIA and ArXivTection datasets to construct a fine-tuning dataset comprising a mix of domains. Then, we fine-tune the LLaMA-7B model on the constructed dataset and evaluate our method on WikiMIA and ArXivTection datasets.
Our results in Table 16 show that our method can also significantly improve the performance of baselines, indicating the effectiveness of our methods when fine-tuning with non-members from a mix of domains. We also evaluate our methods on ArXivTection while fine-tuning using non-members from WikiMIA. The results indicate that our method fails to improve the performance of baselines, since the fine-tuning data comes from an entirely unrelated domain to the evaluation data.
Table 16: AUC score for pretraining data detection with baselines and our method on WikiMIA and ArXivTection under the LLaMA-7B. Wiki (Mix) denote evaluating on WikiMIA and fine-tuning using data from a mix of domains. ArXiv (Wiki) denote evaluating on ArXivTection and fine-tuning on WikiMIA. Base and +Ours respectively denote the baseline methods and our method.
Method Wiki (Mix) ArXiv (Mix) ArXiv (Wiki)
Base +Ours Base +Ours Base +Ours
Perplexity 0.64 0.91 0.68 0.93 0.68 0.52
Lowercase 0.58 0.73 0.50 0.73 0.50 0.50
Zlib 0.62 0.91 0.57 0.92 0.57 0.64
Min-k0.65 0.84 0.76 0.87 0.76 0.61
Table 17: AUC score of baselines and our method on WikiMIA under the LLaMA-7B with different fine-tuning parameters. Base and +Ours respectively denote the baseline methods and our method. Bold shows the superior result.
Method Learning Rate LoRA Rank Epoch
Base $\mathbf{10^{-5}}$ $\mathbf{10^{-4}}$ $\mathbf{10^{-3}}$ Base 8 16 32 Base 1 2 3
Perplexity 0.64 0.81 0.84 0.92 0.64 0.92 0.92 0.92 0.64 0.91 0.91 0.92
Lowercase 0.58 0.60 0.64 0.69 0.58 0.69 0.68 0.69 0.58 0.65 0.64 0.69
Zlib 0.62 0.73 0.78 0.90 0.62 0.91 0.90 0.90 0.62 0.87 0.87 0.90
Min-k0.65 0.76 0.81 0.85 0.65 0.87 0.85 0.86 0.65 0.86 0.87 0.86
(a) AUC
(b) TPR@5%FPR
How do the fine-tuning parameters affect the performance of our method?
To investigate the impact of different fine-tuning parameters on the performance of our method, we conduct experiments on the WikiMIA dataset with different fine-tuning parameters, including learning rate (e.g. 1e-3, 1e-4, 1e-5), epoch (e.g. 1, 2, 3) and LoRA rank (e.g. 8, 16, 32). In Table 17, the results show that our method is relatively insensitive to LoRA rank and the number of fine-tuning epochs. However, considering the learning rate parameter, a learning rate of 0.001 enables our method to perform better.